OpenStudy (anonymous):

sqrt (x+2)+sqrt(x)=4

OpenStudy (anonymous):

Did you not already ask this?

OpenStudy (anonymous):

yea i still dont understand it

OpenStudy (anonymous):

If you square both sides you get \[(\sqrt{x+2}+\sqrt{x})^2=4\]

OpenStudy (anonymous):

Expand and solve.

OpenStudy (anonymous):

x=4-sqrt(x+2)

OpenStudy (anonymous):

all i get is that

OpenStudy (anonymous):

(x+2) + 2sqrt ( x(x+2)) +x = 16

OpenStudy (anonymous):

2x + 2sqrt[x(x+2)] = 14 square again etc